please solve this sequence and find the 7th term.
1,2,6,42,1806..

Answer by scott8148(5895)

1(1 + 1) = 2

2(2 + 1) = 6

6(6 + 1) = 42

42(42 + 1) = 1806

1806(1806 + 1) = 3263442

3263442(3263442 + 1) = 10650056950806

A(n)[A(n) + 1] = A(n+1)